# LLM-based Discriminative Reasoning for Knowledge Graph Question Answering

### v1

@@ 4.2 Main results
The performance of LDR on WebQSP and CWQ is presented in Table 1. The LDR method achieves state-of-the-art performance on these two benchmarks, with improvements in both Hits@1 and F1, indicating an enhanced capability of the LLM to handle KGQA tasks. Despite foregoing the possibility for the LLM to use its own knowledge to directly answer questions correctly, LDR still outperforms previous generative methods, which demonstrates the effectiveness of the proposed strategies (refer to Appendix D for details).
4.3 Enhancement of Subgraph Quality
@@
LDR 0.097 0.209 0.315 0.625 0.181 0.338 1.069 1.490 1.521 3.658 3.000 1.235 1.550 1.578 1.204
Table 2: We use relation recall and minimum graph edit distance as the metrics to measure the quality of retrieved subgraphs with different type of structures. Our method consistently achieves higher relation recall and lower average edit distance than RoG using generation-based searching strategy.
To validate LDR’s enhancement on the capability of LLM to retrieve question-related subgraphs, we design two metrics, relation recall and minimum graph edit distance, to measure the difference between the retrieved subgraph $G_{k}$ and the golden subgraph $G_{gold}$ extract from the SPARQL query given by the benchmarks.
Relation recall measures the proportion of golden relations edges that are successfully predicted, which reflects the method’s sensitivity to retrieve the question-relevant relations in the subgraphs:
$R_{rel}=\frac{count(\{R|R\in G_{k}\}\cap\{R|R\in G_{gold}\})}{count(\{R|R\in G_{gold}\})}.$ (10)
@@
When measuring $D(G_{k},G_{gold})$, the lower the distance, the smaller the structural difference between the predicted graph structure and the correct reasoning subgraph is. We combine the WebQSP and CWQ datasets and classify the test set based on the structure of the given golden subgraph with two features: number of leaf nodes and the total number of relations. The detailed statistic result can be found in Appendix E. We compare LDR with finetuned generative method RoG to evaluate the method’s performance to retrieve different types of subgraphs, the results are shown in Table 2.
Across all types of subgraph structures, it is evident that our method consistently achieves higher relation recall and lower average edit distance, which demonstrates significant enhancement of the LLM’s capability to retrieve question-related subgraphs brought by our proposed discriminative searching strategy.
Subtask Strategy WebQSP CWQ
Hits@1 F1 Graph Hits Hits@1 F1 Graph Hits
- LDR 0.840 0.845 0.884 0.802 0.820 0.894
Answer Inference Generate answer 0.761 0.744 0.884 0.684 0.679 0.894
Subgraph pruning Remove entity type 0.764 0.776 0.884 0.741 0.770 0.894
Ignore constraints 0.737 0.764 0.884 0.548 0.632 0.894
Subgraph Searching Chain-like searching 0.739 0.803 0.896 0.444 0.581 0.844
Table 3: Ablation study of the strategies to remormulate KGQA process as discriminative searching and reasoning process. Those strategies each corresponds to one of the LDR subtasks abbreviate as subgraph retrieval, subgraph pruning and answer inference.
4.4 Reduction of Redundant Information
Figure 3: The average size (number of triplets) of retrieved subgraphs.
Following the analysis of subgraph quality, we calculated the average size of the retrieved subgraphs. As shown in Figure 3, there is a significant reduction in the average size of the retrieved subgraphs, indicating that the LDR method effectively improves the efficiency of subgraph retrieval by recalling fewer but higher-quality subgraph triples.
@@ 4.4 Reduction of Redundant Information
Redundant information may significantly increase the context length, thereby affecting the performance of the generative reasoning process. We analyze the impact of the size of the retrieved subgraph (i.e., the number of triples included) on the overall performance of the strategy. In addition to using RoG with finetuned chain-of-thought reasoning, we implement the in-context reasoning strategy proposed by ToG with the subgraphs retrieved by LDR. To ensure fairness, we use Llama2_7b as the base model for all experiments. As shown in Figure 4, as the number of recalled subgraph triples increases, the performance of generative reasoning methods declines, whereas the strategy adopted by LDR remains more stable. This observation suggests that generation-based reasoning strategies are more sensitive to the size of the subgraphs compared to the discriminative strategy employed by LDR.
4.5 Mitigation of Reasoning Hallucinations
@@
Figure 5: Case frequency of different types of errors.
@@
To analyze the effect of adopting discriminative strategies on ungrounded reasoning with hallucinatory behavior, we collected and examined the frequency of error cases in LDR. Since we decompose the KGQA process into three subtasks executed sequentially, we can categorize all errors into the following three types: 1) E1—failed subgraph searching; 2) E2—insufficient subgraph pruning; 3) E3—incorrect answer location (refer to Appendix D for more details). We analyzed the frequency of these different error types, and the results are shown in Figure 5.
@@
According to the result, compared to the generation-based method, LDR significantly reduces the frequency of E3 errors on both benchmarks and also reduces the overall frequency of E2 errors, which proves that LDR alleviates ungrounded reasoning behaviors in most cases by applying discriminative reasoning strategies. The dropped case frequency of E1 is consistent with our claim that the LDR enhances subgraph searching capability of the LLM.
4.6 Ablation Study of LDR
To validate the effectiveness of the discriminative strategies adopted in LDR, we ablate the strategies we adopted for three subtasks in LDR and observe the changes in performance: 1) Generate answer: ablate our discriminative answer inference strategy and let the model generate the answer based on subgraph $G_{k}$; 2) Remove entity type: replace semantic nodes in $S_{k}$ with unified names "entity_id"; 3) Ignore constraints: skip the subgraph pruning process; 4) Chain-like searching: forbid branching in the subgraph searching subtask, which is similar to the strategy used in RoG. We use "Graph Hits" to measure the proportion of retrieved subgraphs containing at least one correct answer:
$Graph\_Hits=\begin{cases}1,ifA\cap G_{k}\neq\emptyset\\ 0,ifA\cap G_{k}=\emptyset\end{cases}$ (12)
The results are shown in Table 3, which validates the effectiveness of all the LDR strategies.
@@ 4.6 Ablation Study of LDR
Table 4: We change the base model and report the performances. Replacing base models does not have significant impact on the performance of LDR.
0010101010252525255050505075757575100100100100000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.8proportion of training data (%)ReasonSearch
(a) WebQSP
(b) CWQ
4.7 Further Analysis
Model Universality of LDR. To analyze the model Universality of LDR, we test the performance of the LDR method based on different backbone models, and the results are shown in Table 4. The results indicate that changing the base model has no significant impact on the method’s performance, highlighting the universality of the LDR approach.
Utilization Efficiency of Training Data. We examine the utilization efficiency of training data in each LDR subtask, with the results shown in Figure 6. We combine subgraph pruning and answer inference into a single reasoning component and fine-tune two separate models from scratch using Llama2-7b. When evaluating the performance of one model, we use the other model in its fully fine-tuned form. Both models require only about 50% of the training data to achieve performances close to those models trained with entire dataset. Moreover, in terms of data requirements among subgraph retrieval, subgraph pruning and answer inference, subgraph retrieval demands more data and poses greater challenges to the LLM.
Case Study. We present cases of solving KGQA problems using the LDR method in Appendix G. LDR provides effective explicit intermediate reasoning information, which adds to the explainability of the overall KGQA process.
5 Related Works
One of the challenges in the research of KGQA task is to determine how to enable models to learn reasoning processes over knowledge graphs ( Zhu et al., 2024). Early approaches attempted to use networks such as key-value memory networks or graph neural networks to learn representations of inference paths Miller et al. (2016); Yasunaga et al. (2021)). Other works tried to teach models to construct database queries for knowledge graphs (such as SPARQL), allowing them to directly retrieve answers from the graph Gu and Su (2022); Ye et al. (2022)). However, with the advent of large language models (LLMs), the focus has shifted towards leveraging the reasoning capabilities of these models for knowledge inference. For example, UniKGQA ( Jiang et al., 2022) unifies the graph retrieval and reasoning process into a single model using LLMs.
Recently, a popular trend has been the design of effective interactive tools that enable models to engage with knowledge graphs to complete the reasoning process step-by-step, as seen in approaches like KGAgent (Jiang et al., 2024). LLMs have demonstrated surprising capabilities in long-horizon planning and reasoning (Zhong et al., 2024; Wang et al., 2024). By designing a structured interaction process, ToG has achieved remarkable zero-shot reasoning abilities on several KGQA benchmarks with the assistance of GPT-4 (Sun et al., 2024a). FAME (Hong et al., 2023) introduces Monte Carlo planning to generate reliable reasoning steps, while KD-CoT (Wang et al., 2023a) retrieves relevant knowledge from KGs to produce faithful reasoning plans for LLMs. RoG employs chain-like subgraph planning and distills GPT-4’s CoT (Chain-of-Thought) reasoning capability to achieve reliable reasoning processes over knowledge graphs, achieving state-of-the-art performance (LUO et al., 2024).
@@ 6 Conclusion
In this paper, we propose a novel LLM-based discriminative reasoning method, LDR, to address the existing issues of generation-based KGQA methods. Experimental results show that by adopting discriminative strategies, the LDR method not only enhances the capability of LLMs to retrieve question-related subgraphs but also alleviates ungrounded reasoning caused by hallucinations in the generation process.
@@
In the future, we will focus on the following directions to further extend the current work: 1) Conduct analysis and research on KGQA reasoning behaviors and existing issues using scaled models, including GPT-4; 2) Although this work has improved interaction efficiency compared to previously proposed interactive methods (see Appendix H), exploring efficient interaction methods between large models and knowledge graphs remains a valuable research direction; 3) Research on explainable KGQA reasoning processes at the step level: Based on a clear reasoning process, future work can further develop KGQA methods with fine-grained explanations at the level of operational steps, enhancing the explainability and reliability of the KGQA process.
@@ Appendix A Semantic Entity Types
Appendix B Dataset preprocessing details
Due to the huge volume of established Freebase knowledge graph, directly interacting with Freebase through SPARQL is inefficient and may result in unnecessary syntax errors. Following UniKGQA (Jiang et al., 2022), we extract subgraphs from Freebase using breadth-first search for each question, which are then used for the subgraph searching process. Additionally, we expand these subgraphs using the SPARQL queries provided in the benchmarks to ensure the presence of constraint branches.
@@ Appendix B Dataset preprocessing details
In previous work, subgraphs obtained using shortest path search methods typically formed chain-like structures. Compared to the information contained in SPARQL, these structures: 1) might not be logically coherent search paths, and 2) could miss some branches on certain nodes along the path. To enable our method to proceed smoothly, we extracted additional subgraph structures with all possible branches related to the question from Freebase based on the structural information inherent in the SPARQL queries. These were added to the original dataset (for a reference to the original dataset, see RoG). The specific implementation logic can be found in the corresponding functions in the open-source code, and will not be elaborated here.
@@ Appendix C Metrics
Hits@1 Hits@1 calculates the proportion of questions for which the first answer given by the model is correct. Given $A_{pre}$ is the predicted list of answers, and $A_{gold}$ is the list of golden answers, $A_{pre}[0]$ as the very first answer the model predict, then we have:
@@
With this incorrect calculation, the Hits@1 would be higher. For the above example, the Hits@1 for this question would incorrectly become 1.
Recall, precision and F1 We adopt the same calculation method as previous work, using the Macro-F1 scoring method. First, we calculate the precision and recall for each test sample. Then, we average them based on the number of samples to obtain the overall recall and precision. Finally, we use the harmonic mean of the overall recall and precision to calculate the overall F1 score.
Appendix D Types of Errors

### v2

@@ 5.2 Main results
The performance of READS on WebQSP and CWQ is presented in Table 1. According to the results, our porposed READS framework achieves state-of-the-art performance on these two benchmarks, with improvements in both Hits@1 and F1, indicating an enhanced capability of the LLM to handle KGQA tasks. Besides, the READS method abandons the use of internal model knowledge, yet still achieves better KGQA performance, which sufficiently demonstrates that the proposed framework can effectively enhance the knowledge reasoning capabilities of LLMs (refer to Appendix D). We also test READS on more challenging dataset GrailQA Gu et al. (2021), the results are shown in Appendix E.
5.3 Searching Capability Analysis
@@
READS 0.097 0.209 0.315 0.625 0.181 0.338 1.069 1.490 1.521 3.658 3.000 1.235 1.550 1.578 1.204
Table 2: We use relation recall and minimum graph edit distance as the metrics to measure the quality of retrieved subgraphs with different type of structures.
To validate READS’s enhancement on the capability of LLM to retrieve question-related subgraphs, we design two metrics, relation recall and minimum graph edit distance, to measure the difference between the retrieved subgraph $G_{k}$ and the golden subgraph $G_{gold}$ extract from the SPARQL query given by the benchmarks.
Relation recall measures the proportion of golden relations edges that are successfully predicted, which reflects the method’s sensitivity to retrieve the most relevant relations towards the given question from the knowledge graph:
$R_{rel}=\frac{\text{count}(\{R|R\in G_{k}\}\cap\{R|R\in G_{gold}\})}{\text{count}(\{R|R\in G_{gold}\})}.$ (9)
@@
The lower the distance $D(G_{k},G_{gold})$ is, the smaller the structural difference between the predicted graph structure and the correct reasoning subgraph is. We combine the WebQSP and CWQ datasets and classify the test set based on the structure of the given golden subgraph with two features: number of leaf nodes and the total number of relations. The detailed statistic result can be found in Appendix G. We compare READS with finetuned generative method RoG to evaluate the method’s performance to retrieve different types of subgraphs, the results are shown in Table 2.
Across all types of subgraph structures, it is evident that our method consistently achieves higher relation recall and lower average edit distance, which demonstrates significant enhancement of the LLM’s capability to search for question-related subgraphs with our proposed searching strategy.
5.4 Pruning-Answering Capability Analysis
Following the analysis of searching capability, we move on to evaluate the pruning and answering capability of our proposed strategies. We calculate the average size of the retrieved subgraphs. As shown in Figure 3, there is a significant reduction in the average size of the retrieved subgraphs, indicating that the READS method effectively improves the efficiency of subgraph retrieval by recalling fewer but higher-quality subgraph triples.
[…]
(a) WebQSP
[…]
(b) CWQ
Figure 3: The number of cases with the size (number of triplets) of retrieved subgraphs.
@@ 5.4 Pruning-Answering Capability Analysis
We analyze the impact of the size of the retrieved subgraph (i.e., the number of triples included) on the overall performance of the strategy (the result is shown in Figure 4). In addition to using RoG with finetuned chain-of-thought reasoning, we implement the in-context reasoning strategy proposed by ToG with the subgraphs retrieved by READS. To ensure fairness, we use Llama2_7b as the base model for all experiments. As shown in Figure 4, as the number of recalled subgraph triples increases, the performance of generative reasoning methods declines, whereas the strategy adopted by READS remains more stable. Redundant subgraph significantly increase the context length, thereby affecting the performance of the generative reasoning process. This observation suggests that generation-based reasoning strategies are more sensitive to the size of the subgraphs compared to the strategy employed by READS.
5.5 Subtask Ablation Study
To validate the effectiveness of the reformulation approach adopted in READS, we ablate the strategies in READS one at a time and observe the changes in performance. The implementations are: 1) answer: Generate the answer based on the subgraph $G_{k}$ rather than determining position on $S_{k}$; 2) pruning: Skip the pruning process and rely on answer generation process to filter answers; 3) searching: Directly generate subgraph paths based on the question using the strategy in RoG. 4) entity type: Erase the entity type on $S_{k}$, an extra ablation implementation to evaluate the effectiveness of entity information retention for LLM reasoning.
Model WebQSP CWQ
Hits@1 F1 Hits@1 F1
READS 0.840 0.845 0.802 0.820
- answer 0.761 0.744 0.684 0.679
- pruning 0.737 0.764 0.548 0.632
- searching 0.739 0.803 0.444 0.581
- entity type 0.764 0.776 0.741 0.770
Table 3: Ablation study of the strategies in READS.
The results are shown in Table 3. Firstly, all three tasks experienced a performance decline when employing strategies similar to previous work, demonstrating the effectiveness of the task framework proposed by READS. Secondly, the subgraph search task showed the greatest performance difference before and after ablation, indicating that the model’s subgraph search capability is the most critical within the current framework. Lastly, entity type information has been proven to effectively assist large models in conducting more precise reasoning processes.
5.6 Error Type Analysis
@@
Figure 5: Case frequency of different types of errors, E1 corresponds to searching subtask; E2 corresponds to pruning subtask; E3 corresponds to answering subtask.
@@
To analyze the effect of adopting READS on ungrounded reasoning with hallucinatory behavior, we collected and examined the frequency of error cases in READS. Since we decompose the KGQA process into three subtasks executed sequentially, we can categorize all errors into the following three types: 1) E1 stands for abscense of answer in the retrieved subgraph, corresponding to searching subtask; 2) E2 stands for lack of filering of the answer set, corresponding to pruning subtask; 3) E3 stands for mistakenly chosen the position of answer, corresponding to answering subtask (refer to Appendix F for more details). We analyzed the frequency of these different error types, and the results are shown in Figure 5.
@@
According to the result, compared to the generation-based method, READS significantly reduces the frequency of E3 errors on both benchmarks and also reduces the overall frequency of E2 errors, which proves that READS alleviates ungrounded reasoning behaviors in most cases by applying discriminative reasoning strategies. The dropped case frequency of E1 is consistent with our claim that the READS enhances subgraph searching capability of the LLM.
5.7 Hallucination vs. Internal Knowledge
As mentioned earlier, to mitigate hallucinations, READS abandons the capability of LLMs to generate answers directly, focusing instead on discriminative subtasks. To further analyze the balance between hallucination and LLM internal knowledge, we conduct experiments with a strong model GPT-4 and our base model llama2-7b using different answering strategies.
We employed three strategies: 1) Zero-shot: directly listing answers based on the question; 2) READS: using our proposed framework; 3) Augmented: generating answers based on subgraphs extracted by READS. Since we could not fine-tune or constrain the generation process of GPT-4, we presented it with a pool of options and asked it to make a selection.
Strategy WebQSP CWQ
GPT-4 Llama2 GPT-4 Llama2
Zero-shot 0.632 0.403 0.483 0.297
READS 0.544 0.840 0.346 0.802
Augmented 0.856 0.791 0.792 0.632
Table 4: The Hits@1 under different strategies.
As shown in Table 4, compared to the READS process, GPT-4 achieves better results than Llama2 in generating answers based on subgraphs. We believe this illustrates the differences in strategic adaptability between scaled models and 7b-size models. For the commonly used 7b-size models, applying a constrained generation framework may better enhance their ability to inference answers.
5.8 Subtask Data Efficiency
We examine the training efficiency of the READS subtasks. We combine subgraph pruning and answer inference in the figure as reasoning component and fine-tune two separate models from scratch using Llama2-7b. When evaluating the performance of one model, we use the other model in its fully fine-tuned form.
0010101010252525255050505075757575100100100100000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.8proportion of training data (%)Hits@1ReasonSearch
(a) WebQSP
(b) CWQ
As shown in Figure 6, both models require only about 25% of the training data to reach near the best performances. In terms of data requirements among subgraph retrieval, subgraph pruning and answer inference, subgraph retrieval demands more data and poses greater challenges to the LLM.
5.9 Further Analysis
Model Universality of READS. To analyze the model Universality of READS, we test the performance of the READS method based on different backbone models, and the results are shown in Table 5. The results indicate that changing the base model has no significant impact on the method’s performance, highlighting the universality of the READS approach.
@@ 5.9 Further Analysis
Table 5: Model universality of READS.
Reasoning Cost. The interactive analysis between LLMs and knowledge graphs can be quite time-consuming, particularly when large subgraphs introduce long contexts that further hinder reasoning efficiency. However, through the implementation of a highly efficient reasoning strategy, READS has significantly reduced both the average number of model calls per question and the number of tokens per request. As demonstrated in Table 6, READS has halved the cost and achieved a similar average number of model calls as RoG, which plans a subgraph through a single generation.
Method WebQSP CWQ
input output calls input output calls
RoG 343.3 47.4 4.0 490.1 42.9 4.0
ToG - - 11.2 - - 14.3
READS 178.9 10.8 3.9 206.6 12.4 5.7
Table 6: The average model calls per question and average number of input/output tokens per request.
Case Study. We present cases of solving KGQA problems using the READS method in Appendix I. READS provides effective explicit intermediate reasoning information, which adds to the readability of the overall KGQA process.
@@ 6 Conclusion
In this paper, we propose a novel LLM-based reasoning framework READS to reformulate KGQA process, aiming to alleviate the hallucination issues in existing generative methods and enhance the LLM’s reasoning capability. Experimental results proves our claim that by decomposing KGQA and adopting designed discriminative strategies, we can enhances the capability of LLMs to retrieve question-related subgraphs and mitigate ungrounded reasoning results caused by hallucinations in the generation process.
@@
Though our proposed READS framework has shown competitive KGQA performance and is proven to enhance the LLM’s reasoning capability, we identify several limitations that requires further improvement. In the future, we will focus on the following directions to extend the current work:
1) Entity linking: Existing methods assume that the entity linking process is done before the KGQA process LUO et al. (2024); Sun et al. (2024a); In this work we follow the previous works to assume that the entity linking has already been completed. This is a common issue faced by the KGQA methods, we will explore how to eliminate this assumption to achieve reliable KGQA process.
2) Demand on labeled data: Although our method effectively enhances the knowledge reasoning capabilities of large models and demonstrates competitive performance across multiple datasets, we assume the existence of a gold query. Given the strong zero-shot KGQA capability and reasoning capability of GPT-4, works that does not rely on a gold query either requires GPT-4 to annotate the reasoning process (such as RoG) or combines the knowledge memory of strong models to improve overall performance (such as ToG, GoG, etc.) LUO et al. (2024); Sun et al. (2024a); Xu et al. (2024). In the future works, we will explore the possibility of using model-generated pseudo-labels or constructing self-summarized memories to deal with this issue.
@@ Appendix A Semantic Entity Types
Appendix B Data Preprocessing
Training Data for Subgraph Searching. We make use of the SPARQL data available in existing benchmarks to form the training data. In WebQSP and CWQ, each question is associated with a SPARQL query. The direct execution of this query yields the answer to the open question. We obtain the correct subgraph structure required to solve each problem by decomposing the SPARQL statements. Unlike ROG (LUO et al., 2024), in finetuning process READS always presents the model with the correct knowledge subgraph structure rather than the shortest path starts from the question entity and ends at the answer entities.
Training Data for Subgraph pruning and answer inference. To finetune the LLMs to be capable of constraint determination and answer inference, we also construct constraint/answer locating samples from the SPARQL queries in WebQSP and CWQ. The input is a complete subgraph structure with all feasible options of constraints or answer positions, the golden output is the correct position of the constraint and the answer.
Freebase preprocessing. Due to the huge volume of established Freebase knowledge graph, directly interacting with Freebase through SPARQL is inefficient and may result in unnecessary syntax errors. Following UniKGQA (Jiang et al., 2022), we extract subgraphs from Freebase using breadth-first search for each question, which are then used for the subgraph searching process. Additionally, we expand these subgraphs using the SPARQL queries provided in the benchmarks to ensure the presence of constraint branches.
@@ Appendix B Data Preprocessing
In previous works, subgraphs obtained using shortest path search methods typically formed chain-like structures. Compared to the information contained in SPARQL, these structures: 1) might not be logically coherent search paths, and 2) could miss some branches on certain nodes along the path. To enable our method to proceed smoothly, we extracted additional subgraph structures with all possible branches related to the question from Freebase based on the structural information inherent in the SPARQL queries. These were added to the original dataset (for a reference to the original dataset, see RoG). The specific implementation can be found in the corresponding functions in the open-source code, and will not be elaborated here.
@@ Appendix C Metrics
Hits@1. Hits@1 calculates the proportion of questions for which the first answer given by the model is correct. Given $A_{pre}$ is the predicted list of answers, and $A_{gold}$ is the list of golden answers, $A_{pre}[0]$ as the very first answer the model predict, then we have:
@@
With this incorrect calculation, the Hits@1 would be higher. For the above example, the Hits@1 for this question would be 1.
F1. We adopt the same calculation method as previous work, using the Macro-F1 scoring method. First, we calculate the precision and recall for each test sample. Then, we average them based on the number of samples to obtain the overall recall and precision. Finally, we use the harmonic mean of the overall recall and precision to calculate the overall F1 score.
Leaf Node Number 2 3 4 5
with threshold 0.847 0.741 0.625 0.604
w/o threshold 0.845 0.750 0.636 0.738
Method GrailQA Dev
i.i.d compositional zero-shot overall
PANGU 0.844 0.746 0.716 0.754
READS 0.921 0.759 0.626 0.718
Table 9: The Hits@1 performance on GrailQA.
Appendix D Ungrounded Reasoning Behavior
The previous generation-based method can sometimes provide the correct answer even when the subgraph does not contain the correct answer, whereas READS does not exhibit this behavior (see case C2 in Table 11).
Appendix E Result on GrailQA
We test our proposed READ on more challenging benchmark GrailQA’s development set Gu et al. (2021), the results are shown in Table 9. Compared to the previous single task discriminative method PANGU Gu et al. (2023), although not achieving overall SOTA performance, READS enhanced the KGQA reliability on both i.i.d and compositional questions, which proves the effectiveness of the reformulation strategy used in READS.
Appendix F Types of Errors